OpenStudy (anonymous):

13+4a+2a=3a=1 i combined ed like terms and i got 13+9a=3a+1 whats next??

OpenStudy (anonymous):

13 + 4a + 2a = 3a + 1 ---- is this your problem...is the symbol in front of the 1 supposed to be a plus sign ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so, then... 13 + 4a + 2a = 3a + 1 -- combine like terms 13 + 6a = 3a + 1 -- subtract 3a from both sides 13 + 6a - 3a = 3a - 3a + 1 -- combine like terms 13 + 3a = 1 -- subtract 13 from both sides 13 - 13 + 3a = 1 - 13 -- combine like terms 3a = - 12 -- divide both sides by 3 3a/3 = -12/3 a = - 4 any questions ?

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

how would you do this one -8n-4=-2+n-6-8n